# An Improved Two-Step Attack on CRYSTALS-Kyber

### v1

@@ Abstract.
After three rounds of post-quantum cryptography (PQC) strict evaluations conducted by the national institute of standards and technology (NIST), CRYSTALS-Kyber has successfully been selected and drafted for standardization from the mid of 2022. It becomes urgent to further evaluate Kyber’s physical security for the upcoming deployment phase. In this paper, we present an improved two-step attack on Kyber to quickly recover the full secret key, $\mathbf{s}$, by using much fewer energy traces and less time. In the first step, we use the correlation power analysis (CPA) attack to obtain a portion of guess values of $\mathbf{s}$ with a small number of energy traces. The CPA attack is enhanced by utilizing both the Pearson and Kendall’s rank correlation coefficients and modifying the leakage model to improve the accuracy. In the second step, we adopt the lattice attack to recover s based on the results of CPA. The success rate is largely built up by constructing a trail-and-error method. We implement the proposed attack for the reference implementation of Kyber512 (4 128-value groups of s) on ARM Cortex-M4 and successfully recover a 128-value group of s in about $9$ minutes using a $16$-core machine. Additionally, in that case, we only cost at most $60$ CPA guess values for a group and $15$ power traces for a guess.
@@ 1. Introduction
Traditional public-key cryptography Rivest-Shamir-Adleman (RSA) algorithm (rivest1978method, ) and elliptic-curve cryptography (ECC) (koblitz1987elliptic, ) rely on the computational intractability of integer decomposition and discrete logarithm problems respectively. However, concerns have been raised with the emergence of quantum computing because they can be cracked in polynomial time by Shor’s algorithm (shor1994algorithms, ), and thereby revealing the security of existing cryptographic algorithms is insufficient. Recognizing this problem, the national institute of standards and technology (NIST) started the post-quantum cryptography (PQC) standardization process with the aim of standardizing quantum-resistant cryptographic algorithms in 2016 (avanzi2017crystals, ). By July 2022, NIST released the post-quantum cryptographic standard candidates including three signatures and a key encapsulation mechanism (KEM) algorithm in the third round (alagic2022status, ). CRYSTALS-Kyber (bos2018crystals, ) is that only KEM.
Kyber is a lattice-based cryptographic algorithm constructed based on the module-learning with errors (M-LWE) problem. Even in quantum computing, the M-LWE problem is considered to be secure (avanzi2019crystals, ). It should be noted that the mathematical security of Kyber has been widely recognized by the cryptography community. However, for the upcoming deployment phase, it becomes urgent to emphasize its physical security.
Kocher et al. first introduced side-channel attacks (SCAs) in 1996 by leveraging the data dependency on power consumption of cryptographic devices (kocher1996timing, ). Generally, in the absence of any protective measures, devices running the cryptographic algorithms with long-term keys are susceptible to SCAs such as simple power analysis (SPA) attack, template attack (TA), and correlation power analysis (CPA) attack (avanzi2019crystals, ). Different kinds of SCAs for Kyber have gradually emerged. In terms of SPA, Xu et al. successfully conducted such attack on the inverse number theoretic transform (INTT) of the reference implementation of Kyber512 and the pqm4 implementation by constructing specific ciphertext pairs, recovering a coefficient of a secret key costs $8\sim 960$ power traces (xu2021magnifying, ). For TA, Works in (chari2003template, ; choudary2014efficient, ; mu2022voltage, ) analyzed power traces collected from a large number of devices using belief propagation techniques to construct attack templates, enabling the recovery of keys from individual power traces. All of them require tremendous extra data to establish good templates. For CPA, most of previous works (ravi2020generic, ; ueno2022curse, ; shen2023find, ) are chosen-ciphertext attacks (CCAs). The latest work in (shen2023find, ) can significantly reduce the power traces compared to prior works by using an efficient two-step scheme to deal with the imperfect SCA oracles. However, since the NIST PQC KEMs are CCA secure, all of them need extra efforts such as plaintext checking to verify the final results. In (yang2023chosen, ), Yang et al. tried to carry out a random ciphertext CPA attack on the reference implementation of Kyber512, successfully recovering two coefficients of secret key using $20$ or more power traces within a few minutes. Meanwhile, they have demonstrated that the chosen ciphertext CPA attack can improve the efficiency in some degree. Nevertheless, They adopted the original CPA method and the rest of coefficients of the secret key are needed to be recovered separately in the same way. In (kuo2023lattice, ), Yen-Ting Kuo and Atsushi Takayasu recently presented a novel two-step attack on Kyber by integrating the random ciphertext CPA attack with the lattice attack. They constructed a lower dimension of M-LWE problem in the NTT process based on the guess values of CPA and directly calculated the full key. Two hundreds simulated traces are used in their experiments to recover the secret key with about 20 minutes on a 16-core machine for Kyber512. However, they only conducted computer simulations. In this paper, we take the solution of (kuo2023lattice, ) as the starting point and develop an improved two-step attack method on Kyber to further validate it in practical and improve the efficiency.
The proposed two steps can be overviewed as Fig. 1. In this first step, we apply an enhanced CPA attack to recover parts of coefficients of secret key by exploiting the combined correlation between the modified Hamming weight (HW) for some intermediate values and the power consumption of the decryption process in Kyber, specifically the point-wise multiplication covering a secret polynomial and a ciphertext. In this way, some of the secret coefficients can be recovered using a small amount of power traces in number theoretic transform (NTT) domain. In the second step, We take the points of the lattice attack used in (kuo2023lattice, ) and construct a trail-and-error algorithm to recover the entire secret key. Our main contributions can be summarized as follows:
  • We use both Pearson and Kendall’s rank correlation coefficients (Kendall’s tau) and modify the leakage model to improve the accuracy of CPA attacks.
  • Based on lattice attack, we construct a trail-and-error algorithm to improve the success rate.
  • We combine the two steps of attacks together and apply to Kyber512 on ARM Cortex-M4. Experimental results show that the proposed method only costs $38\sim 60$ CPA guess values and about $9$ minutes on a $16$-core machine to recover the 128 coefficients of a secret key s, much faster than the state-of-the-art.
The rest of this paper is structured as follows. Section 2 provides preliminary content. Section 3 presents the traditional attacks and our attack method. Experimental results are described in Section 4. Section 5 summarizes the paper.
Figure 1. The flowchart of the proposed attack.
@@ 2. Preliminaries
In this section, we will introduce the notations, the principles of the LWE/M-LWE problem, NTT in Kyber, CPA Attack, and Kendall’s tau.
2.1. Notations
The ring of integers modulo the prime number $q$ is denoted as $\mathbb{Z}_{q}$. Each polynomial ring $\mathcal{R}_{q}$ = $\mathbb{Z}_{q}[x]/(x^{n}+1)$ with moduli $x^{n}+1$ and $q$ has $n$ coefficients. The Greek symbols $\beta_{\eta}$ and $\mathcal{U}$ denote the centered binomial distribution with parameter $\eta$ and the uniform distribution, respectively. The vectors are represented by bold lowercase letters, such as $\mathbf{a}$, and the vectors in the NTT domain have a hat added to them, such as $\hat{\mathbf{a}}$. The matrices are represented by bold uppercase letters, such as $\mathbf{A}$.
Table 1. Parameter in CRYSTALS-Kyber
$k$ $n$ $q$ $d_{u}$ $d_{v}$ $\eta_{1}$ $\eta_{2}$
Kyber512 2 256 3329 10 4 3 2
Kyber768 3 256 3329 10 4 2 2
Kyber1024 4 256 3329 11 5 2 2
2.2. LWE and M-LWE
Regev et al. introduced the LWE problem (regev2009lattices, ), which forms the basis for several NIST PQC candidates. As show in (1), the LWE problem involves recovering the invariant secret vector $\mathbf{s}$ from $m$ equations, where fixed $\mathbf{s}\leftarrow\beta_{\eta}(\mathbb{Z}_{q}^{n})$.
(1) $(\mathbf{a}_{i},b_{i}=\mathbf{a}_{i}^{\top}\mathbf{s}+e_{i})\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q},$
where $\mathbf{a}_{i}\leftarrow\mathcal{U}(\mathbb{Z}_{q}^{n})$ and the error vector $e_{i}\leftarrow\beta_{\eta}(\mathbb{Z}_{q})$. The essence of the M-LWE problem lies in replacing the ring $\mathbb{Z}_{q}$ in the above LWE problem with the polynomial ring $\mathcal{R}_{q}$, and the error distribution is $\beta_{\eta}(\mathcal{R}_{q})$. Thus, an M-LWE sample $(\mathbf{a}_{i},b_{i})\leftarrow\mathcal{U}(\mathcal{R}_{q}^{k\times 1}\times\mathcal{R}_{q})$ can be represented as:
(2) $(\mathbf{a}_{i},b_{i}=\mathbf{a}_{i}^{\top}\mathbf{s}+e_{i})\in\mathcal{R}_{q}^{k\times 1}\times\mathcal{R}_{q},$
where $\mathbf{a}_{i}\leftarrow\mathcal{U}(\mathcal{R}_{q}^{k\times 1})$ and the error vector $e_{i}\leftarrow\beta_{\eta}(\mathcal{R}_{q})$. Thus, a set of m M-LWE samples can be integrated as:
(3) $(\mathbf{A},\mathbf{b}=\mathbf{A}\mathbf{s}+\mathbf{e})\in\mathcal{R}_{q}^{m\times k}\times\mathcal{R}_{q}^{m}.$
@@ 2.3. NTT in CRYSTALS-Kyber
Kyber is the only standardized KEM of PQC algorithms established by NIST in the third round (moody2021nist, ) and is built on the M-LWE problem. It offers three NIST security levels: Kyber512 corresponding to Level 1, Kyber768 to Level 3, and Kyber1024 to Level 5. The specific parameters are shown in Table 1.
The public-key encryption (PKE) scheme involved in KEM of Kyber consists of three stages: key generation, encryption, and decryption. In the key generation stage, the public key $pk$ is constructed as $\mathbf{A}\mathbf{s}+\mathbf{e}$, where $\mathbf{A}$ is the sampling polynomial matrix. $\mathbf{s}$ and $\mathbf{e}$ are the secret key and noise, respectively, both of which are polynomial vectors sampled from the $\beta_{\eta_{1}}$. During the encryption stage, the message $m$ is encrypted into ciphertext $c$, where $c$ is formed by compressing a polynomial vector $u$ and an array $v$ and concatenating them. In the decryption stage, the receiver extracts $u$ and $v$ from the ciphertext $c$, and then utilizes $\mathbf{s}$ to perform corresponding operations to recovery the message $m$. Algorithm 1 illustrates the decryption process of Kyber. The KEM protocol is an extension of the PKE protocol with re-encryption. The application of the Fujisaki-Okamoto transformation (fujisaki1999enhance, ) to an IND-CPA-secure PKE results in an IND-CCA2-secure KEM.
@@
In Kyber, polynomial multiplication is a fundamental operation that is frequently used in the encryption and decryption processes. By performing multiplication calculations on the polynomial converted to the NTT domain, the polynomial multiplication on NTT can reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$, thus accelerating the speed of the entire encryption and decryption process. Note that for Kyber, here is only $256^{\text{th}}$ primitive root of unity $\xi$ instead of $512_{\text{th}}$. Therefore, the modulus $x^{n}+1$ in Kyber can only be partially factored into $n/2$ quadratic polynomials, with odd and even coefficients calculated respectively. It is also considered that NTT is a linear transformation, the specific formula is described as follows:
[…]
@@
where $br$($i$) represents a 7-bit bit-reversal of $i$ and $\mathbf{M}_{n\times n}$ is a reduced integer matrix. Similarly, INTT can be represented in the same way.
Algorithm 1 Kyber.CPAPKE.Dec(sk,c):decryption
1: Input: Secret key $sk\in\mathcal{B}^{12\cdot k\cdot n/8}$
2: Input: Ciphertext $c=(\mathbf{u},v)\in\mathcal{B}^{d_{u}\cdot k\cdot n/8+d_{v}\cdot n/8}$
3: Output: Message $m\in\mathcal{B}^{32}$
4:$\mathbf{u}:=\text{Decompress}_{q}(\text{Decode}_{d_{u}}(c),d_{u})$
5:$v:=\text{Decompress}_{q}(\text{Decode}_{d_{v}}(c+d_{u}\cdot k\cdot n/8),d_{v})$
[…]
8:return $m$
2.4. Correlation Power Analysis Attack
SCAs on cryptographic devices can be broadly categorized into invasive, non-invasive, and semi-invasive attacks. Non-invasive attacks, such as the CPA attack, can compromise secret keys without disrupting the operation of the cryptographic devices. In the following, we will delve into the principles and steps involved in CPA attack in summary.
In practical SCAs, physical phenomena such as the power consumption and the electromagnetic radiation are often observed (verbauwhede2010secure, ). The CPA attack is a widely used method in SCAs, which relies on the correlation between power models such as HW or Hamming distance (HD) and actual power consumption (joy2011side, ). The higher the correlation guess, the greater the likelihood that the key with higher correlation guess is the correct one. A traditional CPA attack can generally be divided into four steps.
Firstly, select an intermediate value $\mathit{f}(d,k)$ computed by the device’s encryption algorithm as the attack point, where $d$ is a known partial ciphertext or message and $k$ is a partial key. Secondly, measure the actual power consumption of the device. When the device encrypts or decrypts $D$ different messages or ciphertexts, we record the actual power consumption $\mathbf{T}_{D\times T}$, where $\mathbf{T}_{ij}$ denotes the power consumption of the $i^{th}$ plaintext or ciphertext at time $j$. Thirdly, calculate the hypothetical intermediate values and map them to the real power consumption. The hypothesis intermediate $\mathbf{V}=\left[f(d_{i},gk_{j})\right]_{D\times K}$ is computed for all guess keys $gk$ and mapped to the power consumption $\mathbf{H}_{D\times K}$. Finally, calculate the Pearson correlation coefficient (PCC) between $\mathbf{T}$ and $\mathbf{H}$ to obtain the correlation coefficient matrix $\mathbf{P}_{K\times T}$. According to the largest value in $\left|\mathbf{P}\right|$, we determine the correct part of the private key $k_{c}$ and time. The full secret key can be recovered by repeating the above procedure for each partial secret key.
2.5. Kendall’s Rank Correlation Coefficient
The Kendall’s tau is a non-parametric measure of the association between two random variables. It quantifies the degree of concordance in the rankings of two variables, i.e., whether they are consistently ranked in terms of their values (abdi2007kendall, ). Kendall’s tau $\tau$ ranges from $-1$ to $1$, where $\tau=1$ indicates a perfect direct association, $\tau=-1$ indicates a perfect disassociation, and $\tau=0$ indicates no agreement in rankings between the two variables.
[…]
where $c$ and $d$ denote the numbers of concordant and discordant pairs, while $t_{x}$ and $t_{y}$ represent the counts of tied ranks in the $x$ and $y$ data sets, respectively.
@@ 3. Proposed Improved Two-Step Attack
It should be remarked that the proposed improved two-step attack is illustrated in Fig. 1. We will detail the basis of the attack analysis and our proposed three improvements in the following.
3.1. Basis of Attack Analysis
The goal of this paper is to explore the physical security of Kyber KEM under random-ciphertext CPA attack. The purpose of the attack is to obtain a long-term secret key. We assume that an attacker has access to a device which is running Kyber decryption and can enter arbitrary ciphertext into the device. In addition, they can also capture the electromagnetic radiation of the device to obtain the power traces.
In subsection 2.3, we have introduced NTT used in Kyber briefly, which only possesses $n_{\text{th}}$ roots of unity because the modulus polynomial $(x^{n}+1)$ can only be factored into $n/2$ linear polynomials. Consequently, multiplication involves the multiplication of linear polynomials rather than simple point-wise multiplication in the NTT domain. Due to the properties of incomplete NTT in Kyber, the full secret key of Kyber can be divided into $2k(k=2,3,4)$ groups.
The test vector leakage assessment (TVLA) methodology can be used to find side-channel leakage points and the t-test method is one of the commonly used methods in TVLA (wu2022efficient, ). Several t-test results from previous works have indicated the potential leakage points during the decryption process in Kyber (yang2023chosen, ; cryptoeprint:2022/058, ). One critical side-channel leakage occurs during the point-wise multiplication of $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$. Additionally, this operation takes place in the quotient ring $\mathbb{Z}_{q}[x]/(x^{2}-\xi)$. In the clean implementation of pqm4 (kannwischer2019pqm4, ), polynomial multiplication employs 256 $basemuls$ in Kyber512. Moreover, each calculation of $basemul$ is independent from the others.
From Listing 1, we can observe that there are two steps related to the partial secret key coefficient $s_{0}$ and three steps related to the partial secret key coefficient $s_{1}$. The inputs of a $basemul$ are $s_{0}$, $s_{1}$, $u_{0}$, and $u_{1}$, and outputs are $r_{0}$ and $r_{1}$. Five multiplications and two additions are involved in the listing. Obviously, the result is related to two secret key coefficients. All partial secret key coefficients range from $\left(0,3328\right)$ for Kyber. In a practical CPA attack, suppose that the attacker aims to recover the coefficient $s_{0}$, $r_{0}$ can be selected as the intermediate value. Similarly, if the attacker aims to recover the coefficient $s_{1}$, $r_{1}$ can be selected as the intermediate value. In addition, in order to recover the full secret key coefficients of a group, the attacker would need to perform 128 CPA attacks. Therefore, $128\times 2k$ CPA attacks are required for the full secret key.
Listing 2: Modified HW in the CPA Attack.
1def HW(r0):
2 if r0 < 0:
3 r0 = calculate_complement(r0, 16)
4 H_W = r0.count("1")
5 else:
6 H_W = bin(r0).count("1")
7 return H_W
8
9def calculate_complement(r0, 16):

### v2

# An Improved Two-Step Attack on Kyber

@@ Abstract.
After three rounds of post-quantum cryptography (PQC) strict evaluations conducted by the national institute of standards and technology (NIST), CRYSTALS-Kyber was successfully selected in July 2022 and standardized in August 2024. It becomes urgent to further evaluate Kyber’s physical security for the upcoming deployment phase. In this brief, we present an improved two-step attack on Kyber to quickly recover the full secret key, $\mathbf{s}$, by using much fewer energy traces and less time. In the first step, we use the correlation power analysis (CPA) to obtain a portion of guess values of $\mathbf{s}$ with a small number of energy traces. The CPA is enhanced by utilizing both the Pearson and Kendall’s rank correlation coefficients and modifying the leakage model to improve the accuracy. In the second step, we adopt the lattice attack to recover s based on the results of CPA. The success rate is largely built up by constructing a trial-and-error method. We implement the proposed attack for the reference implementation of Kyber512 (4 128-value groups of s) on ARM Cortex-M4 and successfully recover a 128-value group of s in about $9$ minutes using a $16$-core machine. Additionally, in that case, we only cost at most $60$ CPA guess values for a group and 15 power traces for a guess.
@@ 1. Introduction
WITH the fast development of quantum computing, modern public-key cryptographic algorithms are encountering significant threats. Recognizing this problem, the national institute of standards and technology (NIST) started the post-quantum cryptography (PQC) standardization process in 2016 with the aim of standardizing quantum-resistant cryptographic algorithms. By July 2022, NIST released the post-quantum cryptographic standard candidates including three signatures and one key encapsulation mechanism (KEM) algorithm after the third round evaluation (alagic2022status, ). CRYSTALS-Kyber is that only KEM, which has been successfully standardized as module-lattice-based KEM (ML-KEM) noted as FIPS 203 from August 2024 (Announcing2024, ). It should be marked that the mathematical security of Kyber has been widely recognized by the cryptography community. However, for the upcoming deployment phase, it becomes urgent to emphasize its physical security (nagata2021physical, ; zhao2023side, ; ravi2024side, ).
Kocher et al. first introduced side-channel attacks (SCAs) in 1996 by leveraging the data dependency on power consumption of cryptographic devices (kocher1996timing, ). Generally, in the absence of any protective measures, devices running the cryptographic algorithms with long-term keys are susceptible to SCAs, especially to the correlation power analysis (CPA).
Recent years, several CPA attacks on Kyber (ravi2020generic, ; ueno2022curse, ; shen2023find, ; yang2023chosen, ; kuo2023lattice, ) have been proposed by researchers. In 2020, Ravi et al (ravi2020generic, ). used plaintext-checking (PC) oracle-based SCAs to recover the secret key of Kyber512 with 2560 power traces according to the vulnerabilities in Fujisaki-Okamoto (FO) transform for the first time. Then, Ueno et al (ueno2022curse, ). proposed the side channel leakage in re-encryption as a PC oracle, utilizing 3072 power traces to recover the key. The work in (shen2023find, ) deals with imperfect PC oracle SCAs by using the CPA attack twice to recover 60% secret key of Kyber512 at the cost of about 1619 power traces, but there are still 40% error coefficients. Contrast to previous works, in (yang2023chosen, ), Yang et al. tried to carry out a random ciphertext CPA attack on the reference implementation of Kyber512, successfully recovering two coefficients of secret key using $20$ or more power traces. Nevertheless, they had to repeat the attack 256 times to recover the entire secret key of Kyber512. So, they at least need 5120 power traces for a key. In (kuo2023lattice, ), Yen-Ting Kuo and Atsushi Takayasu presented a two-step attack on Kyber by integrating the random ciphertext CPA attack with the lattice attack. They adopted the CPA attack to recover part of the coefficients of a key and then directly solved the rest coefficients by using a low-dimension equation of M-LWE problem. The correctness was verified by using computer simulations. In this paper, we take the idea of (kuo2023lattice, ) as the starting point and develop an improved two-step attack method on Kyber to further validate it in practical and improve the efficiency.
Figure 1. The flowchart of the two-step attack.
The proposed two-step attack is summarized in Fig. 1. In the first step, we modify the leakage model to better align our CPA with the arithmetic logic unit. Additionally, we introduce an optional correlation coefficient, the Kendall rank correlation, to address the nonlinearity and robustness issues of the original CPA in real-world attacks. In the second step, we construct a trial-and-error lattice attack algorithm to recover the entire secret key, and optimize the parameters of the algorithm. By employing this two-step attack, we achieve significant improvements in both efficiency and accuracy in practice.
@@ 2.1. NTT in CRYSTALS-Kyber
Computing polynomial multiplication in NTT domain can reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$, thus significantly improving the speed of the entire encryption and decryption process in Kyber. Note that for Kyber, here is only $256^{\text{th}}$ primitive root of unity $\xi$ instead of $512^{\text{th}}$. Therefore, 256 coefficients are divided into two groups, namely the odd and even groups, for NTT calculations. The even coefficients in NTT domian are described as follows:
(1) $\hat{f}_{2i}=\sum_{j=0}^{127}f_{2j}\xi^{(2br(i)+1)j}\Rightarrow\ [\hat{f}_{2i}]^{\top}=\mathbf{M}[f_{2i}]^{\top},0\leq i<128,$
@@
where $br$($i$) represents a 7-bit bit-reversal of $i$ and $\mathbf{M}$ is a $128\times 128$ matrix with reduced integers. Similarly, INTT can be represented in the same way and the inverse matrix is denoted as $N$.
@@
2.2. Correlation-Based Attack Techniques
The conventional CPA attack is a non-invasive side-channel attack that uses the Person correlation between power consumption models and actual power traces to recover cryptographic keys. It identifies the correct key by finding the highest correlation between measured and hypothetical values. Pearson correlation coefficient (PCC) is efficient and well-suited for detecting linear relationships between the hypothetical and actual values.
In contrast, Kendall’s tau is a non-parametric measure of association that quantifies the degree of concordance between the rankings of two variables, indicating whether their values are consistently ranked together (abdi2007kendall, ). Kendall’s tau $\tau$ ranges from $-1$ to $1$, with $\tau=1$ indicating perfect agreement, $\tau=-1$ indicating perfect disagreement, and $\tau=0$ indicating no agreement. The Kendall’s tau calculation is computed as follows:
(2) $\tau_{b}=\frac{c-d}{\sqrt{({c+d+t_{x}})({c+d+t_{y}})}},$
where $c$ and $d$ denote the numbers of concordant and discordant pairs, $t_{x}$ and $t_{y}$ represent the counts of tied ranks in the $x$ and $y$ data sets, respectively.
We have separately implemented the two methods in software and found that the PCC computation is much faster than the Kendall’s tau calculation. The main reason is that the latter must compute a large full permutation. Luckily, when we combined the two methods together, the efficiency and reliability both were enhanced. More details will be shown in the following section.
@@ 3. Proposed Improved Two-Step Attack
It should be remarked that the proposed improved two-step attack is illustrated in Fig. 1. We will detail the basis of the attack analysis and our proposed three improvements in the following. We assume that an attacker has access to a device which is running Kyber decryption and can only observe ciphertexts entered into the device. In addition, they can also capture the power traces of the device.
Several test vector leakage assessment (TVLA) results from previous works have indicated the potential leakage points during the decryption process in Kyber (shen2023find, ; yang2023chosen, ). One critical side-channel leakage occurs during the point-wise multiplication of $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$. In the clean implementation of pqm4 (kannwischer2019pqm4, ), polynomial multiplication employs 256 $basemuls$ in Kyber512. Moreover, each calculation of $basemul$ is independent from the others. From Listing 1, it is clear that the result depends on two secret key coefficients, $s_{0}$ and $s_{1}$. All partial secret key coefficients range from $\left(0,3328\right)$ for Kyber. In a practical attack, suppose that the attacker aims to recover the coefficient $s_{0}$. $r_{0}$ can be selected as the intermediate value.